Market demand in the duration of 20 days follows normal distribution. The demand data for
20 days follow the series {100, 110, 120,…}. Consider the random variable X which takes the
values X=100, 110, 120,… Find the expectation of X and also variance of X
1
Expert's answer
2020-04-15T10:43:38-0400

"E(X)=\sum XP(X)"

X follows an arithmetic series with "a=100" and "d=10"

Sum of 20 terms is given by "10(200+(10\u00d719))=3900"

"P(X)=\frac{1}{20}"

Thus, "E(X)=\frac{3900}{20}=195"

"Var(X)=E(X^2)-(E(X))^2"

"X^2" Will be 10000, 12100, 14400...

Thus, "\sum X^2=827000"

"E(X^2)=\frac{82700}{20}=41350"

"Var(X)=41350-195^2=3325"